OpenStudy (anonymous):

How to simplify this? \( \frac{-\sqrt{3} - 1}{1+(-\sqrt{3}*1}\)

OpenStudy (anonymous):

It is suppose to be How to simplify this? \( \frac{-\sqrt{3} - 1}{1+(-\sqrt{3})*1}\)

OpenStudy (gentorr):

3 11 11 3

OpenStudy (anonymous):

HUH? the answer is \( 2+ \sqrt{3}\) but how did they get it?

TheSmartOne (thesmartone):

First simplify the denominator.

OpenStudy (anonymous):

How?

TheSmartOne (thesmartone):

\(\sf\Large 1+ (-\sqrt{3})\times 1\) Anything times 1 is 1 and when you add a negative number you are basically subtracting it. Make sure to follow PEMDAS

OpenStudy (anonymous):

So I need to How to simplify this? \( \frac{-\sqrt{3} - 1}{1+(-\sqrt{3}*1} \frac{1+(-\sqrt{3}*1}{1+(-\sqrt{3}*1}\)

TheSmartOne (thesmartone):

We are getting to that, first we need to simplify the denominator :)

OpenStudy (anonymous):

So \[ \frac{-\sqrt{3} - 1}{1+(-\sqrt{3}*1} \frac{1+(-\sqrt{3})}{1+(-\sqrt{3})} \]

OpenStudy (anonymous):

\[ \frac{-\sqrt{3} - 1}{1+(-\sqrt{3}} * \frac{1+(-\sqrt{3})}{1+(-\sqrt{3})} \]

TheSmartOne (thesmartone):

yes, however: \(\sf 1+(-\sqrt{3})=1-\sqrt{3}\)

TheSmartOne (thesmartone):

and then we have to multiply the numerator and denominator by the conjugate of \(\sf1-\sqrt{3}\)

OpenStudy (anonymous):

Do we use foil to times \[ (-\sqrt{3}-1)(1-\sqrt{3}) \]

TheSmartOne (thesmartone):

yes

OpenStudy (anonymous):

I got 2 for the numerator is this correct?

TheSmartOne (thesmartone):

I fail to see how you got that...